Grade 12th passGeneral Physics

A ball is thrown vertically upward with a speed of 120m/s. It's speed at midpoint of it's path is about ( g=10)

Amit kumar
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear student
 
v^2 = u^2 – 2as
14400 =2 * 10 * s 
 
s = 720 m
 
now
 
360 = 120t – 5 t^2
 
t^2 – 24 t + 72 = 0
 
now take t from here and use this in
v =u + at
 
hope it helps